OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 6) and a directrix at y = -6.

OpenStudy (anonymous):

@bradely

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

vertex = midpoint of directrix and focus = (0, 0)

ganeshie8 (ganeshie8):

So the equation will be of form : \[\large y = \dfrac{1}{4c}x^2\]

ganeshie8 (ganeshie8):

c=6 here, cuz directrix is y = -6 = -c that makes the required equation : \[\large y = \dfrac{1}{4*6}x^2\]

ganeshie8 (ganeshie8):

\[\large y = \dfrac{1}{24}x^2\]